Circular No. 6013

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Word has been received of the independent discoveries of a comet by

Masamitsu Nakamura, Hideo Nishimura, and Donald Machholz. The following positions are available:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 5.713  3^{h}58.^{m}2  +70^{o}11′  9.0 Nakamura
5.749  3 58  +70 10  9  Nishimura
6.408  3 56  +70 12  10.5 Machholz
6.43663 3 55 26.^{s}28 +70 12 25′′.5 10  Helin
6.46528 3 55 21.56 +70 12 26.4  _{”}

M. Nakamura (Hamamatsu, Shizuoka-Ken, Japan). 25×150 binoculars.

Coma diameter 5′. Communicated by T. Seki and S. Nakano. H. Nishimura (Kakegawa, Shizuoka-Ken, Japan). 25×150 binoculars. Com-

municated by S. Nakano. Coma diameter 2′.5. D. Machholz (Colfax, CA). 27×120 binoculars. Diffuse with some conden-

sation. E. Helin, K. Lawrence, and J. Alu (Palomar). 0.46-m Schmidt telescope

films. Measured by Lawrence. Comet is diffuse with strong central condensation.

COMET TAKAMIZAWA-LEVY (1994f) Total visual magnitude estimates (cf. IAUC 6006): June 8.90 UT, 8.8

(K. Sárneczky, Ráktanya, Hungary, 20×60 binoculars); 14.04, 8.8 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); 28.27, 8.7 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 30.93, 9.6 (A. Diepvens, Balen, Belgium, 0.15-m refractor); July 2.29, 9.4 (Spratt); 3.29, 9.7 (Spratt).

NOVA OPHIUCHI 1994 A. C. Gilmore, Mount John University Observatory, reports the fol-

lowing additional photometry (cf. IAUC 6008; uncertainty ∼ 0.02 in each color except U − B, where the error may be 0.05): June 21.617 UT, V = 10.68, U − B = −0.41, B − V = +0.48, V − R_{c} = +1.25, V − I_{c} = +1.34.

1994 July 6 Daniel W. E. Green